40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1327     Entry time: Thu Feb 19 23:50:31 2009
Author: pete 
Type: Update 
Category: Locking 
Subject: aligned pd's on AP table 

Yoichi, Peter

While continuing our efforts to lock, we noticed the procedure failed at a point it had gotten past last night:  turning on the bounce/roll filters in MICH, PRC, and SRC.  We checked the MICH transfer function and noticed that the unity gain point was ~10 Hz, well below the bounce modes.   We tried increasing the gain but found saturation, and Rob suggested that there could be misalignment on the AP table, which Steve worked on today.  We went out and found two of the PDs (ASDD133 and AS166) to be badly misaligned probably due to a bumped optic upstream.  We re-aligned.

 

 

ELOG V3.1.3-